### v2

@@ Abstract
Patient management requires multitasking interaction with multimodal data. While today’s artificial intelligence (AI), particularly large foundation models, promises unprecedented opportunities, progress remains relatively slow in developing medical multimodal multitask foundation models. There are two main challenges along this direction: the data challenge – the high bar to curate medical multimodal multitask datasets including 3D or higher dimensional medical tomographic images in alignment with other clinical datasets, and the model challenge – the unavailability of a scalable and adaptable foundation model architecture to synergize multimodal datasets for diverse clinical tasks. Here we propose the first-of-its-kind medical multimodal-multitask foundation model (M3FM) with an emphasis on chest CT tasks especially lung cancer screening. To train our M3FM, we first curated a comprehensive multimodal multitask dataset consisting of 163,725 3D chest CT exams, 48 clinical data types, and 17 medical tasks on lung, heart, and other chest diseases. Then, we created and applied a multimodal question-answering framework as a unified training strategy to effectively integrate multimodal information and naturally perform multiple tasks with free-text prompting. Our extensive experimental results demonstrate that M3FM consistently outperforms the previous state-of-the-art single-modality task-specific models. Also, M3FM can identify informative multimodal data elements that are relevant to specific clinical tasks, being instrumental in building AI models and gaining insights into correlations among the multimodal data and diseases. Furthermore, M3FM can be adapted to boost the performance of new tasks with a small out-of-distribution dataset. Importantly, M3FM can handle different combinations of multimodal datasets some of which could be incomplete and manage high-dimensional medical images at different scales depending on tasks. Our M3FM has enabled superior volumetric CT imaging performance for lung cancer screening, cardiac disease prediction, and other CT-related tasks. This foundation model can be extended to incorporate more data types and improve other medical tasks, towards AI-empowered precise and efficient medicine.
@@ 1 Introduction
Modern medical diagnosis and treatment require simultaneous multiple tasks on multimodal information with imaging, structured laboratory data including genomics, and unstructured free-text data from radiology, pathology, and clinical notes. Taking lung cancer screening (LCS) as an example, radiologists need to detect and characterize lung nodules, as well as identify lung, heart, and other chest abnormalities in medical images of a patient, and summarize all findings into a radiology report. Then, the patient will be managed by taking all the results and risk factors into account. In the LCS process, multimodal data are involved, including chest X-ray, low-dose computed tomography (LDCT) images, patient demographics, smoking history, disease history, family cancer history, pathological results, follow-up data, etc. [1]. The optimization of LCS outcomes demands great expertise of radiologists and other physicians, honed through years of training and experience [2]. The involved workload for radiologists is significant, incurs substantial costs, and hinders a broader implementation of high-quality personalized medicine that can improve population health.
@@
For LCS, it has been demonstrated that volumetric LDCT screening reduced lung cancer mortality by 20% in comparison with chest radiography in the National Lung Screening Trial (NLST) [1] and reduced 24% in comparison with no screening in the NELSON trial [3]. A follow-up US study [4] reported that the USA states with a lower LCS fraction among eligible adults bear a higher lung cancer burden. However, the LCS fraction of eligible smokers remains low relative to other screening tests (such as for cervical, colorectal, and breast cancers) due to various barriers to implementing LCS, such as financial constraints and resource inaccessibility [5]. The situation becomes more challenging with the increasing number of lung cancer patients who have no or little smoking history [6, 7] and the sub-optimal management for LCS [8, 9, 10]. Additionally, a global shortage of radiologists with the requisite training for providing LCS poses a significant barrier to its widespread implementation and accessibility for patients. Hence, there is an important and immediate need for multidisciplinary efforts to broadly, equitably, and optimally implement LCS for reduced lung cancer mortality [5].
Artificial intelligence (AI) models are increasingly recognized as powerful tools that align with the necessity for transitioning to value-based healthcare by improving quality of clinical services while reducing human workload and overall cost. Compared to human intelligence (HI), AI models are beginning to provide analyses and predictions beyond the capabilities of human experts capabilities in various clinical scenarios. For example, deep learning methods were proposed for lung cancer detection and risk estimation with LDCT in an end-to-end manner [11]. Recently, the Sybil model [12] was developed for lung cancer risk prediction using a single LDCT scan. Since LCS using LDCT also offers an opportunity for assessing cardiovascular disease (CVD) risks [13], a deep learning model was developed for CVD risk prediction with LDCT, achieving promising results on both NLST and independent evaluation datasets [14]. All these deep learning methods are focused on narrow tasks using single modality data, which do not take multimodal data into systematic consideration. Due to the high dimensionality of volumetric CT images, these methods used small 2D/3D ResNet models [15] and down-sampled 3D LDCT images so that the computational cost is affordable. In particular, the training schemes of current lung cancer risk models [11, 12] require costly bounding box annotations, which makes it impractical to leverage large-scale datasets to scale up advanced AI models.
The latest generation AI methods are large foundation models(FMs), which are usually built with large-scale transformers [16] and trained on massive and diverse datasets. Large FMs have updated the state-of-the-art (SoTA) performance across a wide range of downstream tasks [17], such as large language models (LLM) [18, 19, 20], computer vision models [21, 22], vision language models [23, 24], and generalist models [25, 26]. Propelled by breakthroughs in artificial general intelligence (AGI), the medical community calls for generalist medical AI (GMAI) [27]. GMAI models are expected to perform diverse tasks per instructions in free text, making predictions with different combinations of data modalities and interpreting the outputs as medical professionals. Towards this ambitious goal, the transfer of AI successes to healthcare is gaining momentum. Remarkably, some biomedical chatbots have been developed by fine-tuning open-source LLMs [28, 20] on customized biomedical datasets, such as ChatDoctor [29], Med-Alpaca [30], PMC-LLaMA [31], Clinical Camel [32], DoctorGLM [33], and Huatuo [34]. These biomedical chatbots are useful to assist specialists. Recently, the closed-source LLMs (PaLM [35] and Flan-PaLM [36]) were tuned with a curated medical language dataset [37]. By the scaling laws [38], the resulting model, Med-PaLM [37], achieves SoTA performance on various datasets. All these results suggest the potential of LLMs in medicine although they can only process textual data. Using contrastive learning techniques [23, 39], a number of vision-language models [40, 41, 42, 43] have been designed to identify abnormalities in 2D chest X-ray images [44, 45, 46, 47]. In clinical practice, 3D LDCT is significantly more precise for LCS than 2D chest radiography [1] and thus recommended by the US Preventive Services Task Force [48]. However, these models are limited in interpreting volumetric or higher-dimensional medical images along with various clinical data.
@@
It is well recognized that given the sheer size of multimodal datasets covering diverse medical tasks, the larger AI models could yield unprecedentedly better medical AI outcomes [38], thereby revolutionizing numerous clinical scenarios. However, progress remains relatively slow in this direction for two main reasons. First, we face the data challenge. There is a high bar to curate medical multimodal multitask datasets, systematically aligning 3D or high-dimensional medical images and other diverse clinical data with various medical tasks in specific clinical scenarios. Recent studies either combined multiple public datasets covering diverse tasks while each task mainly depends on a single modality data [49] or collected large-scale biomedical figure-caption datasets from PubMed for general image captioning tasks [50]. However, these data collection and curation strategies are not well aligned with specific clinical scenarios. Second, we have the model challenge. There is no scalable and adaptable foundation model dedicated to effectively interpreting medical multimodal data, especially high-dimensional medical images at different scales, and flexibly performing diverse clinical tasks. Although some efforts [49, 50] have explored fine-tuning general domain foundation models for medical tasks, how to effectively encode and interpret multimodal data and how to synergistically unify diverse medical tasks remain an open clinical challenge. As in the GPT-4V(ision) assessment  [51], current models are not sufficiently developed to support real-world medical imaging and decision-making. In particular, their inherited vision encoders are pre-trained with natural images and can only process 2D images of limited size. Therefore, more dedicated multidisciplinary efforts are urgently needed to build clinical-grade FMs that synergize multitasks and multimodal data, particularly including tomographic image volumes.
In this article, we report the first-of-its-kind medical multimodal multitask FM (M3FM) with an emphasis on chest CT tasks, especially lung cancer screening. The success of our M3FM model is achieved by addressing both the above-articulated data and model challenges.
@@
First, we present an integrated and scalable data curation approach to align high-dimensional medical images with other clinical datasets for diverse tasks in an overarching clinical context by leveraging domain-specific expertise. Specifically, we define a set of comprehensive AI tasks for the whole process of LCS, as summarized in Figure 2. To collect the data for the defined tasks, we were granted access to massive data from the National Lung Cancer Screening Trial (NLST) and the Medical Imaging and Data Resource Center (MIDRC), both of which involved multiple medical institutes. Also, we independently collected two similar datasets from Wake Forest University School of Medicine and Massachusetts General Hospital respectively. Our collected datasets take a variety of data types, including chest CT scans and the associated imaging parameters, radiology reports, de-identified demographics, smoking history, disease history, cancer history, family history, pathological results, and follow-up data. We systematically processed and aligned the multimodal data for each defined task, resulting in 64 sub-datasets for training, validating, testing, and fine-tuning, as summarized in Section 2.1. It is worth noting that we used only medical records without expensive human labeling, such as bounding box annotations for lung nodules or other imaging findings, during the data curation so that the dataset can be easily enlarged.
@@
Second, we develop the first-of-its-kind medical multimodal multitask foundation model (M3FM) architecture with an emphasis on LCS with LDCT. M3FM can not only synergistically analyze different combinations of medical multimodal datasets including multi-scale 3D tomographic scans but also simultaneously perform multiple tasks in a unified and scalable manner. We are motivated by the observation that radiologists routinely undertake multiple tasks, drawing upon multimodal information to generate a comprehensive radiology report. Specifically, radiologists repeatedly scroll through CT volume slices multiple times, each time concentrating on a specific task and focusing on target organs/tissues within a particular CT display window, until all tasks of interest are finished. Radiologists also access patient data in the electronic medical record to gain necessary insight regarding relevant medical history that may help refine their insights from the imaging exam. Taken together, this workflow demands that radiologists receive a task command at a time and then only pay attention to the relevant information in a serial fashion, which must be ultimately synthesized for making accurate diagnoses. To mimic this procedure, we formulate our M3FM in a multimodal question-answering (MQA) framework. As shown in Fig. 1, M3FM encodes multimodal data, incorporates prior knowledge, and predicts the answer to each given question. The prior knowledge involves different aspects: the correspondence between organs/tissues and the target task, clinically used display windows, and candidate answers for each task. With prior knowledge, M3FM takes as input the anatomically relevant sub-volume in a specific display window for the given task. By processing relevant sub-volume only, both the computational cost and redundant/irrelevant information can be reduced. However, 3D large organs, such as the lungs, still have a much larger size than common 2D images. Hence, novel memory-efficient techniques are the key to training M3FM on commonly used GPUs. Another computational challenge comes from multitask training, which aims to optimize multiple objectives simultaneously using mixed task-specific datasets. To this end, we have designed a task-distributed training strategy, enabling M3FM scalable to multiple tasks. Since abnormalities may have significantly different sizes and correlate with different clinical data portions, M3FM should accommodate various scales of volumetric images and different combinations of multimodal data. For this purpose, we design a CT Vision Transformer (CTViT) compatible with multiple image sizes as an image encoder of M3FM. In particular, CTViT disentangles the physical size from the image and embeds it as a separate input. This makes M3FM flexibly process different scales of images without losing their physical size information, which is an important factor for specific tasks, as demonstrated in Section 2.3. We describe all clinical data with free text so that any combinations of multimodal data can be easily input into the model. During training, we randomly drop out some clinical data to avoid overfitting, ensuring it works with incomplete multimodal data. Since large-scale self-supervised pre-training has become an essential technique to ensure the performance of large foundation models [52, 18, 53, 54], we adapted the masked image modeling method [53, 54] to pre-train our CTViT image encoder with a large number of volumetric CT scans and used the text Transformer pre-trained via masked language modeling [52] as our text encoder.
Our key contributions are summarized as follows: (1) We build the first-of-its-kind medical multimodal multitask foundation model (M3FM) that effectively encodes different combinations of multimodal data, especially high-dimensional tomographic images at different scales, and flexibly performs multiple tasks via free-text prompting. M3FM sets new benchmarks for a variety of tasks in the LDCT LCS process and beyond. (2) We present a systematic workflow to build a clinical grade foundation model, from clinical multitask definition to multimodal data curation, from radiologists’ reading procedure to the unified M3FM framework, from self-supervised pre-training to multitask learning with high-dimensional multimodal data. Importantly, the whole pipeline is designed with scalability in mind, allowing M3FM to be readily scaled up by integrating more training datasets and undertaking a broader range of clinical tasks. (3) Our extensive experimental results demonstrate that by synergizing multimodal data and multitasks, our M3FM model significantly improves the performance of the counterpart model trained on single-modality data for individual tasks, and the informative data types for each task can be identified through ablation inference or training. Also, M3FM can be adapted to boost the performance of new tasks with a small out-of-distribution dataset. (4) We plan to release our curated medical multimodal multitask datasets (OpenM3Chest), our pre-trained CTViT and M3FM models, and source codes for research use. We aspire that our salable M3FM will serve as a foundational architecture to synergize multimodal data and integrate related medical tasks for superior medical AI performance, deriving more advanced specialist and generalist models.
@@
Figure 1: M3FM inference illustration. M3FM consists of four main components: CTViT, Text Transformer, Task Attention, and Predictors. CTViT encodes volumetric CT images at different scales (indicated by rectangle boxes in different sizes and colors). Text Transformer encodes a combination of clinical data (rectangle text boxes) and questions (callout text boxes) in free text. All image and text tokens are forwarded to the Task Encoder, which extracts task-specific embedding features of the integrated multimodal data. The task-specific Predictors take task-specific embedding features to derive the final answers. These tasks may have arbitrary combinations of inputs, varied sizes of image inputs, and shared/different predictors as indicated by color matching/mismatching respectively.
Figure 2: Dataset construction. The green boxes indicates the OpenM3Chest dataset that will be made publicly available.
Figure 3: Multitask definition and multi-modality data.
Table 1: Details of the Multimodal Question-Answering Datasets.
Table 2: Data composition of the OpenM3Chest datasets. This table shows the number of CT series in each dataset for each task and their distributions over the corresponding candidate answers in brackets.
Figure 4: Comparison results of ChestFM with previous SoTA models in terms of AUC.
Figure 5: Results of different scales of models in terms of AUC and 95% CI.

### v3

@@ Abstract
Modern medical records include a vast amount of multimodal free text clinical data and imaging data from radiology, cardiology, and digital pathology. Fully mining such big data requires multitasking; otherwise, occult but important aspects may be overlooked, adversely affecting clinical management and population healthcare. Despite remarkable successes of AI in individual tasks with single-modal data, the progress in developing generalist medical AI (GMAI) remains relatively slow to combine multimodal data for multitasks because of the dual challenges of data curation and model architecture. The data challenge involves querying and curating multimodal structured and unstructured text, alphanumeric, and especially 3D tomographic scans on an individual patient level for real-time decisions and on a scale to estimate population health statistics. The model challenge demands a scalable and adaptable network architecture to integrate multimodal datasets for diverse clinical tasks. Here we propose the first-of-its-kind medical multimodal-multitask foundation model (M3FM) with the initial application in low-dose CT lung cancer screening and related tasks. After we curated a comprehensive multimodal multitask dataset consisting of 49 clinical data types including 163,725 chest CT series and 17 medical tasks involved in lung cancer screening, we develop a multimodal question-answering framework as a unified training and inference strategy to synergize multimodal information and perform multiple tasks via free-text prompting. Our M3FM consistently and significantly outperforms the state-of-the-art single-modal task-specific models, improving lung cancer risk prediction by up to 20% and cardiovascular disease mortality risk estimation by up to 10%. M3FM identifies multimodal data elements informative for clinical tasks, being instrumental in gaining insights and correlating the multimodal data with relevant diseases. M3FM can flexibly adapt to new tasks with a small out-of-distribution dataset, effectively handle various combinations of multimodal data, and efficiently process high-dimensional images at multiple scales. In a broader sense, as a specialty-oriented generalist medical AI (SOGMAI) model, M3FM innovates lung cancer management and related tasks. This SOGMAI approach paves the way for similar breakthroughs in other areas of medicine, closing the gap between specialists and the generalist.
@@ 1 Introduction
Medical diagnosis, treatment, and management require simultaneous multiple tasks on multimodal data, such as unstructured medical records (for symptoms, comorbidities, risk factors, etc.), structured text data (such as laboratory test results), and imaging data (such as computed tomography (CT) images). In the fast-evolving domain of artificial intelligence (AI), foundation models (FMs) have shown previously unseen abilities to understand diverse data types and execute many tasks in a unified architecture [1]. Large FMs have updated the state-of-the-art (SoTA) performance across a wide range of tasks, such as natural language processing [2, 3, 4], computer vision [5, 6], and vision-language understanding [7, 8]. Propelled by the breakthroughs of FMs, the medical community calls for generalist medical AI (GMAI) [9]. GMAI models are expected to perform diverse tasks per instructions in free text, making predictions with different combinations of data modalities and interpreting the outputs as medical professionals.
Along this direction, biomedical language models [10, 11, 12, 13, 14, 15] were developed by fine-tuning open-source [16, 4] or closed-source [17, 18] large language models (LLMs) on the customized biomedical datasets. Among these models, Med-PaLM [19] achieves SoTA performance on various datasets due to its large model size, which demonstrates the power of scaling laws [20] in the medical domain although they can only process textual data. Most recently, Google Research and Deep Mind prepared a multimodal biomedical benchmark called MultiMedBench and developed the Med-PaLM Multimodal (Med-PaLM M) system [21]. On their test datasets, Med-PaLM M reached performance comparable to or better than the SOTA records. While it represents a milestone proof of concept, Med-PaLM M inherits the vision encoder pre-trained with natural images, only takes 2D images of limited size, handles each task mainly by taking a single modality data without systematic alignment among multiple data modalities, and accumulates relatively disconnected public datasets and medical tasks that are not well aligned with specific clinical scenarios. These facts limit its utility in real-world clinical applications.
Beyond the above proof-of-concept studies, we believe that the efforts toward GMAI should involve the development of specialty-oriented generalist medical AI (SOGMAI) models. Integration of these SOGMAI models will bring us to the holy grail of GMAI. It is underlined that SOGMAI is both general and specific. It is general in scope, as it encompasses a multitude of interconnected tasks and interprets a variety of correlated data modalities within each medical specialty. At the same time, the SOGMAI model is specific in alignment with a medical specialty, making it seamlessly applicable to streamline a specific clinical workflow. Here we introduce our Medical Multimodal Multitask Foundation Model (M3FM) as the first SOGMAI model, expressly designed for lung cancer screening (LCS). LCS involves multiple tasks, such as lung nodule detection and characterization, lung cancer risk evaluation, diagnosis of a set of chest abnormalities, COVID-19 detection, and diagnosis and risk evaluation of cardiovascular diseases (CVD); and multimodal data, including low-dose computed tomography (LDCT) images, patient demographics, smoking history, disease history, family cancer history, pathological results, follow-up data, etc [22].
@@
Indeed, lung cancer remains the leading cause of cancer-related deaths [23]. Treatment at an earlier stage is much more effective and less invasive, reducing the medical cost and social burden. For LCS, 3D LDCT is the method of choice over 2D chest radiography, as LDCT screening reduced lung cancer mortality by 20% in comparison with chest radiography in the National Lung Screening Trial (NLST) [22] and 24% mortality reduction in comparison with no screening in the NELSON trial [24]. However, the LCS fraction of eligible smokers remains low relative to other screening tests (such as for cervical, colorectal, and breast cancers) due to various barriers to implementing LCS, such as resource inaccessibility [25, 26], the challenging patient management [27, 28, 29, 30, 31], and particularly, a global shortage of radiologists with the requisite training for providing LCS. Hence, there is an important and immediate need for multidisciplinary efforts to broadly, equitably, and optimally implement LCS for reduced lung cancer mortality [25].
Over the past years, various task-specific AI methods have been studied on the LCS datasets. For example, a deep learning method was proposed for lung cancer detection and risk estimation with LDCT in an end-to-end manner [32]. Recently, the Sybil model [33] was developed for lung cancer risk prediction using a single LDCT scan. In some studies [34, 35], deep learning models were developed for cardiovascular diseases (CVD) risk prediction with LDCT from LCS. All these models are focused on narrow tasks using single modality data, limiting their performance and utility in LCS. Due to the high dimensionality of volumetric CT images, these efforts only studied small 2D/3D ResNet models [36] with affordable computation costs. In particular, the training schemes of current lung cancer risk models [32, 33] require costly bounding box annotations, which makes building large-scale training datasets impractical.
@@
It is well recognized that given the sheer size of multimodal datasets covering diverse medical tasks, the large FMs could yield unprecedentedly better medical AI outcomes [20], thereby revolutionizing numerous clinical scenarios. However, progress remains relatively slow in this direction for two main reasons. First, we face the data challenge. There is a high bar to curate medical multimodal multitask datasets, systematically aligning 3D medical images and other diverse structured and unstructured text-based clinical data with various medical tasks in specific clinical scenarios. For example, current data collection and curation strategies [37, 38] are not well aligned with specific clinical scenarios. Second, we have the model challenge. There is no scalable and adaptable foundation model dedicated to effectively interpreting medical multimodal data, especially high-dimensional medical images at different scales, and flexibly performing diverse clinical tasks. Although some efforts [37, 38] have explored fine-tuning general domain foundation models for medical tasks, how to effectively encode and interpret multimodal data and how to synergistically unify diverse medical tasks remain an open challenge.
In this study, we report the first-of-its-kind SOGMAI model, M3FM, that can perceive multimodal data to multiple tasks dedicated to LCS, as illustrated in Figure 1. The success of our M3FM for LCS is achieved by addressing both the data and model challenges.
@@
First, we present an integrated and scalable data curation approach to align high-dimensional medical images with other clinical datasets for specialty-oriented tasks by leveraging domain-specific expertise. According to the LCS practice, 17 tasks of interest were defined and 49 different data elements were involved. To collect the multimodal data for these tasks, we were granted access to massive data from the National Lung Cancer Screening Trial (NLST) and the Medical Imaging and Data Resource Center (MIDRC), both of which involved multiple medical institutes. Also, we independently collected two similar datasets from Wake Forest University School of Medicine (WFUSM) and Massachusetts General Hospital (MGH). Our datasets were curated by aligning among a variety of data types for each task, including chest CT scans and the associated imaging parameters, radiology reports, de-identified demographics, smoking history, disease history, cancer history, family history, pathological results, and follow-up data. The resulting 64 sub-datasets for training, validating, testing, and fine-tuning are summarized in Figure 2 and Tabel 1. It is worth noting that we used only medical records without expensive human labeling, such as bounding box annotations for lung nodules or other imaging findings, during the data curation so that these datasets could be easily enlarged.
@@
Second, we develop a scalable and adaptable M3FM architecture with an emphasis on LCS with LDCT. M3FM can not only analyze different combinations of medical multimodal data including multi-scale 3D tomographic scans but also perform multiple tasks in a unified manner, the architecture details are shown in Figure 6. Routinely, radiologists undertake multiple tasks, drawing upon multimodal information to generate a comprehensive radiology report. They repeatedly scroll through CT volume slices multiple times, each time focusing on a specific task and its relevant organs/tissues within a particular CT display window. Radiologists also access the electronic medical record to help refine their insights from the imaging exam. This workflow demands that radiologists receive a task command at a time and pay attention to all relevant data to make accurate diagnoses. To mimic this procedure, we formulate our M3FM in a multimodal question-answering (MQA) framework. As shown in Fig. 1, M3FM encodes multimodal data, incorporates prior knowledge, and predicts the answer to each question. With prior knowledge, M3FM takes as input the anatomically relevant sub-volume in a specific display window for the given task. Since different diseases may be distributed in significantly different sizes in CT and correlated with different clinical data portions, M3FM should accommodate various scales of volumetric images and different combinations of multimodal data. For this purpose, we design a CT Vision Transformer (CTViT) compatible with multiple image sizes as an image encoder of M3FM. CTViT disentangles the physical size from the image and embeds it as a separate input, which makes M3FM flexibly process different scales of images without losing their physical size information, which is an important factor for specific tasks, see Section 2.3. We describe all clinical data with free text so that any combinations of multimodal data can be easily input into the model. Since large-scale self-supervised pre-training has become an essential technique to ensure the performance of large foundation models [39, 2, 40, 41], we adapted the masked image modeling method [40, 41] to pre-train our CTViT image encoder with a large number of volumetric CT scans and used the text Transformer pre-trained via masked language modeling [39] as our text encoder. In our extensive evaluation, we find that by synergizing multimodal data and multitasks, our M3FM significantly improves the performance of the counterpart model trained on single-modality data for individual tasks, and the informative data types for each task can be identified through ablation inference or training. As a foundational model, M3FM can be adapted to boost the performance of new tasks with a small out-of-distribution dataset.
@@
Figure 1: M3FM inference illustration. M3FM consists of four main components: CTViT, Text Transformer, Task Encoder, and Predictors. CTViT encodes volumetric CT images at different scales (indicated by rectangle boxes in different sizes and colors). Text Transformer encodes a combination of clinical data and questions in free text. All image and text tokens are forwarded to the Task Encoder, which extracts task-specific embedding features of the integrated multimodal data. The task-specific Predictors take task-specific embedding features to derive the final answers. These tasks may have arbitrary combinations of inputs including varied sizes of imaging data and shared/different predictors. Different colors in CT, QUESTIONS/TASKS, and ANSWERS differentiate the matches among them. The questions in the black text take the textural inputs only. Questions 17 and 18 are two examples of simulated clinical information retrieval tasks to inspect the ability of clinical data modeling. Question 19 only takes its neighboring clinical text as input.
Figure 2: Dataset construction. (a) The general data construction workflow consists of four steps: medical tasks of interest definition, task-specific multimodal data collection, multimodal data processing and alignment, and multimodal question-answering construction. (b) The data used in this study were collected from two data centers, i.e., NLST and MIDRC, and two medical institutes, i.e., WFUSM and MGH, with the key summarized characteristics, based on which a large 3D CT pretraining dataset and a simulated clinical dataset were constructed. The green boxes indicate the OpenM3Chest dataset that will be made publicly available. (c) Distributions of main training, validation, and test datasets over all tasks are summarized. (d) Distributions of MGH independent evaluation datasets are summarized. (e) Distributions of WFUSM independent evaluation and fine-tuning datasets are summarized.